OpenStudy (anonymous):

Soybean meal is 18% protein; cornmeal is 9% protein. How many lbs of each should be mixed together to get a 360lb mix that is 15% protein? (application of linear equations)

OpenStudy (anonymous):

18%x+9%y=360 * 15%

OpenStudy (anonymous):

^equation1

OpenStudy (anonymous):

x+y=360 equation 2

OpenStudy (anonymous):

s+c=360 (18/100)s + 9/100)c=15*360

OpenStudy (anonymous):

solve them and you'll have your answer (linear systems of equations)

OpenStudy (anonymous):

ty hahd

OpenStudy (anonymous):

240.0 and 120 is the answer

OpenStudy (anonymous):

and you are welcome